OpenStudy (anonymous):

Please help me find the roots of following polynomial equation: e^2x-5(e^2x)+6 = 0

OpenStudy (amistre64):

is that a typo? I dont think there is a '2' in the middle term..

OpenStudy (anonymous):

Its not a typo.Its from our exam paper.

OpenStudy (angela210793):

@amistre64...May you please,please,please see my problem? Please??????? :)

OpenStudy (amistre64):

then combine like terms ..... a = e^2x -4a+6 = 0 a = 6/4 = 3/2 e^2x = 3/2 2x = ln(3/2) x = ln(3/2)/2

OpenStudy (amistre64):

if i can find it ;)

OpenStudy (angela210793):

perhaps it is in my profile...try 2 click once there :)

OpenStudy (anonymous):

I am not getting your solution. Can you please explain?

OpenStudy (amistre64):

not getting as you dont follow it? or the answer is wrong?

OpenStudy (anonymous):

I do not follow it.Its kinda tough.

OpenStudy (anonymous):

pleasse come back later to my question, dont quite understand what you mean

OpenStudy (amistre64):

e^2x can be substituted for becasue it doesnt change in the problem; its an unknown that we determine; so to clean up the equation, make it = a and use 'a' instead. a -5a +6 =0 -4a +6 = 0 -4a = -6 a = -6/-4 = 6/4 = 3/2

OpenStudy (amistre64):

ok :)

OpenStudy (amistre64):

since e^2x = a ; we told it to ... e^2x = 3/2 we solve for x by using logarithmic properties.. ln(e^2x) = ln(3/2) 2x = ln(3/2) x = ln(3/2) ------ 2

OpenStudy (anonymous):

Its involved I guess.I would look into it.Thanks for your help.I appreciate it.

OpenStudy (amistre64):

your welcome But I really think that this could have been a typo; which would make this part easier: a = e^x a^2 -5a +6 = 0 (a-6)(a+1) = 0 a = 6 and a = -1 e^x = a e^x = 6 ; x = ln(6) e^x = -1, which is impossible so disregard the -1 x = ln(6)